Aug 23, 2018 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. BCAL and FCAL: Develop standard fiducial volume cuts for the calorimeters
  2. BCAL: Create non-linearity corrections for MC for gain adjustment. (Tegan)
  3. FCAL: Cross check MC changes by Richard to make sure that gains are not affected in both Geant3 and Geant4 (Collin)
  4. BCAL: Provide Sean with an estimate of the number of pi0s needed to do calibrations (inclusive and exclusive channels, cell calibrations as well as non-linear corrections) (Tegan) Information provided to Sean. Mark thought that Sean decided to skim the first 30 files of each run.

Complete after new launch

  1. From Mike's work: -cos(phi) type dependence in the data and a phi dependence at the downstream end of the bcal. See Sep 21, 2017 Calorimeter Meeting.


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Updates
    1. FCAL
    2. BCAL
  4. Calibrations
  5. Monitoring
  6. Reconstruction and Simulation
  7. NIM articles
  8. Any other business


Attending: Elton, Mark, Tegan, Colleen, Will (JLab); Zisis, Ahmed (UofR); Matt (IU); Christina (Athens); Rupesh, Tolga (FIU)

  1. Announcements
    • Run is delayed another day. Potential shortage of nitrogen to the lab might have a severe impact on the current run, but availability is under review.
  2. Action Items
  3. Run Updates
    1. FCAL
      • All base spares have been used. Another dozen bases will be replace when spares are available.
      • Most of the bases have the HV failure. Failures of other types are mounting. When Rebecca was here they mix-matched parts to obtain working spares. Bases with these failures will be gathered and returned to IU for repair.
      • "Bad" cable fixed (connector issue) and FADCS with bad channels have been replaced.
      • One FADC module with new firmware was installed in the FCAL to test new firmware, but there are issues and further testing is needed before routine use.
    2. BCAL
      • BCAL set for nominal operation.
      • Chillers repaired from the corrosion and erosion in the piping. New piping replaced with stainless steel.
      • Upstream chiller had additional troubles. Two parts were identified as problematic and parts have been ordered. The upstream chiller was replaced with the spare, which was going to be used to cool the CompCal.
  4. Calibrations
    • Calibration Dashboard
    • Tegan has updated the BCAL gain constants for 2018.
    • New calibration was installed in time for the 2018 monitoring launch.
    • The pi0 resolution is comparable to last year when using a direct comparison of a new double-Gaussian + pol fit to data to 2018 and 2017 data.
  5. Monitoring
    • Main goal is to check occupancies and 4 ns time shifts in data
    • Sample plots were sent around to a few persons for comments. They show structures that are not completely understood, so they are not yet ready for shift takers. Mark will return comments this afternoon.
    • Downstream pulser seems to be missing some cycles. The upstream pulser is more stable.
    • Mark: There are ROOTSPY histograms that are not routinely displayed by the shift takers but available to experts. This might be an option for the LED monitoring.
    • Zisis: Is it easy to interchange the upstream and downstream trigger cables and see if the downstream electronic triggering modules has an issue? Elton and Mark: Yes, this could be done and we can run a couple of LED test runs with interchanged cables when the DAQ is available.
  6. Reconstruction and Simulation
    1. SiPM saturation (Elton)
      • The SiPM saturation has been implemented at the hit level in both simulation (mcsmear) and reconstruction. New table was added to ccdb.
      • New branch has been tested by looking at the J/Psi events with Lubomir and shows that the observed non-linearity at high energies is mostly corrected with the new code.
      • Implementation of the new software with require a check of the pi0 gain calibration and new non-linearity correction.
      • A note is being prepared with information about the correction and made ready for implementation. The timing of the implementation will need discussion with Sean and considerations based on production of 2018 data.
  7. NIM articles
  8. Any other business